# Review of Machine Learning Approaches for Diagnostics and Prognostics of Industrial Systems Using Industrial Open Source Data

### v1

@@ 6 Discussion and Prospects
Nevertheless, there are still several pivotal suggestions that merit deeper investigation and exploration by the research community going forward.
(1) Advancements in Model Interpretability, Robustness, and Generalization. The adoption of deep learning in PHM has underscored the need for models that are not only high-performing but also interpretable. Techniques such as advanced data visualization and explainable artificial intelligence methods are emerging as key tools in explaining the outputs of machine learning models. These methods are anticipated to provide industries and academia with clearer insights into the decision-making processes of PHM models, thereby building trust and facilitating more informed decision-making. Alongside interpretability, the robustness and generalization of ML models in diverse and unpredictable industrial settings are paramount. Novel approaches in transfer learning and domain adaptation can be further developed to ensure models are resilient to data variability and operational uncertainties and capable of adapting to new, unseen scenarios. This ensures their applicability in a wide range of PHM applications, driving the field towards more efficient, reliable, and transparent systems.
(2) Development of Integrated ML Toolbox. The creation of an integrated ML toolbox is essential for streamlining research in PHM. By consolidating various machine learning methods into a unified toolbox, researchers can quickly access and utilize existing tools, thus accelerating the pace of research and enhancing problem-solving efficiency. Such a toolbox allows for data analysis and ML model construction in a timely manner when encountering familiar problems, leading to substantial time savings. Moreover, the toolbox is not static; it can be continuously updated with new processing ideas and solutions, fostering an environment conducive to the rapid deployment of ML in PHM. However, this requires a collaborative effort across the PHM community.
@@
(3) Construction of Large Knowledge Library. There have seen increasing applications of machine learning in PHM, but the knowledge and solutions remain fragmented. Therefore, there is a need to establish a comprehensive large knowledge library that can include open-source industrial sensor data, problem-solving approaches, open-source code, model weights, etc. This dynamic knowledge library could evolve over time, becoming more comprehensive and thereby benefiting the broader PHM community. For instance, Lee’s team proposed to construct a industrial large knowledge library including human-interpretable data and structured machine-generated data [38]. The proposed industrial LKL could serve as a domain-specific foundation for training Large Knowledge Models to solve domain-specific problems.
@@
Figure 3: Industrial Large Knowledge Model Framework [38].
@@ 7 Conclusion
In summary, machine learning gradually becomes a cornerstone in PHM, reflecting the big potential for innovative advancements in future PHM development. This paper serves as a valuable resource for both academic and industry professionals in PHM domain, offering a unified ML framework in PHM and a comprehensive overview of the current state-of-the-art machine learning approaches for diagnostics and prognostics of industrial systems using industrial open-source data. Based on two primary research task categories: ”Detection & Diagnosis” and ”Assessment & Prognosis”, we provide a detailed explanation of the problems, tasks, challenges, and relevant machine learning approaches to each competition. Furthermore, we summarize common challenges, including data-related issues and model-related issues, and analyze the solutions to address these challenges. Finally, we prospect several potential directions in the application of data-driven machine learning within PHM, encompassing improving the interpretability, robustness, and generalization of ML models, developing integrated ML toolboxes, constructing a industrial large knowledge library, harnessing the potential of LLMs and ILKMs.

### v2

@@ 6 Prospects
There are still several pivotal suggestions that merit deeper investigation and exploration by the research community going forward.
(1) A Need for Open-Source Multi-Modal Datasets. In the PHM domain, the availability of multi-modal datasets is notably limited. While prior research has leveraged diverse modal information, including vibration, current, or torque signals for diagnosing issues in wind turbine gearboxes, bearings, or other industrial products [45, 114, 31], these datasets are all private. This restriction to some extent hampers the capacity for broad-based development of MMML approaches. To overcome this challenge, there is a need for the PHM community to collaboratively establish and maintain multi-modal industrial datasets, enriched with high-quality data and labels. This would involve the collection, alignment, and annotation of multi-modal data with PHM-centric attributes. Moreover, providing latent representations or pre-trained embeddings, if possible, can also accelerate and efficiently train new MMML models and facilitate knowledge transfer across various PHM tasks, ultimately benefiting the whole PHM community.
(2) Development of Multi-Modal Machine Learning Approaches. Furthermore, the investigation of a broader range of MMML techniques is highly encouraged. On the one hand, for time series data of a single modality, it may be feasible to extract features representative of different modalities from the time series data itself, such as the time domain (origin signal), frequency domain (FFT), PSD, STFT, etc. This approach could lead to the preliminary training of unimodal ML models on each single modality, followed by the exploration of MMML strategies. On the other hand, MMML methodologies necessitate more effective representation learning and information alignment techniques—the former concerning the efficient encoding of single modality data, and the latter focusing on the enhanced analysis and fusion of multimodal information for effective PHM prediction or classification tasks. Although simple concatenation is a common method for data fusion in MMML, emerging fusion techniques, such as attention-based or transformer-based mechanisms [111], deserve further exploration. These advanced methods have the potential to effectively capture implicit feature alignments across modalities and facilitate cross-modal synthesis [71, 119]. Yet, research on MMML within the PHM field remains underexplored, highlighting a significant opportunity for the adaptation of existing methods and the innovation of new approaches to address the unique challenges of PHM.
(3) Further Exploration in Machine Learning Model Interpretability. The adoption of DL in PHM has underscored the need for models that are not only high-performing but also interpretable. Techniques such as advanced data visualization and XAI methods are emerging as key tools in explaining the outputs of ML models [64, 6]. These methods are anticipated to provide industries and academia with clearer insights into the decision-making processes of PHM models, thereby building trust and facilitating more informed decision-making [98]. However, current XAI methods predominantly address the interpretability of models using tabular data, text, and images as inputs, leaving a gap in methods tailored for time series data. Moreover, most of the current interpretability methods are applied to unimodal ML models, and the interpretability of deep MMML models has not been explored. Addressing these gaps presents valuable opportunities for future research, aiming to balance the performance and interpretability of ML models. Such advancements would make ML models more trustworthy and interactive, enhancing their utility for scientists and engineers.
(4) Novel Transfer Learning and Domain Adaptation Techniques Development for Model Robustness, and Generalization. Alongside interpretability, the robustness and generalization of ML models in diverse and unpredictable industrial settings are paramount. Novel approaches in transfer learning and domain adaptation can be further developed to ensure models are resilient to data variability and operational uncertainties and capable of adapting to new, unseen scenarios [8]. This ensures their applicability in a wide range of PHM applications, driving the field towards more efficient, reliable, and transparent systems. Currently, research on transfer learning in the PHM domain predominantly addresses fault diagnosis, with only a few studies exploring prognosis. Looking forward, how to utilize the power of transfer learning for prediction problems is still a critical issue. Additionally, cross-modal transfer learning (CMTL) emerges as a critical area of interest in PHM, aiming to improve the knowledge transfer between distinct domains. Moreover, the challenge of collecting a sufficiently large, labeled dataset is a significant barrier in practical applications. The development of unsupervised and semi-supervised transfer learning techniques may help to address this issue.
@@
(5) Construction of Large Knowledge Library. There have seen increasing applications of ML approaches in PHM, but the knowledge and solutions remain fragmented. Therefore, there is a need to establish a comprehensive large knowledge library that can include open-source industrial sensor data, problem-solving approaches, open-source code, model weights, technical papers, review documents, etc. This dynamic knowledge library could evolve over time, becoming more comprehensive and thereby benefiting the broader PHM community. By consolidating various industrial application information into this knowledge library, researchers can quickly access and utilize existing techniques when encountering familiar problems, thus accelerating the pace of research and enhancing problem-solving efficiency. For instance, Lee’s team proposed to construct a industrial large knowledge library including human-interpretable data and structured machine generated data [59]. The proposed industrial LKL could serve as a domain-specific foundation for training Industrial Large Knowledge Models to solve domain-specific problems.
@@
Figure 6: Industrial Large Knowledge Model Framework [59]
@@ 7 Conclusion
In summary, ML gradually becomes a cornerstone in PHM, reflecting the big potential for innovative advancements in future PHM development. This paper serves as a valuable resource for both academic and industry professionals in the PHM domain, offering a unified ML framework in PHM and a comprehensive overview of the current state-of-the-art ML approaches for diagnostics and prognostics of industrial systems using industrial open-source data. Based on two primary research task categories: ”Detection & Diagnosis” and ”Assessment & Prognosis”, we provide a detailed explanation of the problems, tasks, challenges, and relevant ML approaches to each competition. Furthermore, we summarize common challenges, including data-related issues and model-related issues, and analyze the solutions to address these challenges. Moreover, we evaluate the limitations of these PHM data challenge competitions and suggest future directions for the PHM data challenge competition evolution. Finally, we prospect six potential directions in the application of data-driven ML within PHM, encompassing a need for open-source multi-modal datasets, development of MMML approaches, further exploration of ML model interpretability, improving the robustness, and generalization of ML models, constructing a large knowledge library, and harnessing the potential of LLMs and ILKMs.
Nomenclature
$AI$ Artificial Intelligence
$Bi-LSTM$ Bidirectional Long Short-Term Memory
$CNN$ Convolutional Neural Network
$DANN$ Domain Adversarial Neural Networks
$DL$ Deep Learning
$DTW$ Dynamic Time Warping
$FCM$ Fuzzy C-Means
$FCN$ Fully Convolutional Network
$FIR$ Feature Importance Ranking
$GAN$ Generative Adversarial Networks
$GRU$ Gated Recurrent Units
$ILKM$ Industrial Large Knowledge Model
$KNN$ K-Nearest Neighbors
$LDA$ Linear Discriminant Analysis
$LightGBM$ Light Gradient Boosting Machine
$LIME$ Local Interpretable Model-Agnostic Explanations
$LKL$ Large Knowledge Library
$LLM$ Large Language Model
$LSTM$ Long Short-Term Memory
$ML$ Machine Learning
$MLP$ Multi-Layer Perceptron
$MMD$ Maximum Mean Discrepancy
$PCA$ Principal Component Analysis
$PDF$ Probability Density Function
$PHM$ Prognostics and Health Management
$PLS$ Partial Least Squares
$PSD$ Power Spectral Density
$QA$ Question Answering
$RNN$ Recurrent Neural Network
$RUL$ Remaining Useful Life
$SHAP$ SHapley Additive exPlanations
$SMOTE$ Synthetic Minority Oversampling TEchnique
$SoQ$ Stream-of-Quality
$STFT$ Short-Time Fourier Transform
$SVM$ Support Vector Machine
$SVR$ Support Vector Regression
$TCN$ Temporal Convolutional Network
$XAI$ Explainable Artificial Intelligence
$XGBoost$ Extreme Gradient Boosting
Appendix
Table 9: Task Description and Website Links for PHM Data Challenge Competitions
PHM Data Challenge Competition Task Description URL
2018 PHM NA
Ion Mill Etching System
(1) Diagnose failures
(2) Determine time remaining until next failure
https://phmsociety.org/conference/annual-conference-of-the-phm-soc
iety/annual-conference-of-the-prognostics-and-health-management
-society-2018-b/phm-data-challenge-6/
2019 PHM NA
Fatigue Crack
(1) Estimate crack length https://data.phmsociety.org/2019datachallenge/
2020 PHM EU
Filtration System
(1) Predict RUL of the filtration system http://phmeurope.org/2020/data-challenge-2020
2021 PHM EU
Manufacturing Production Line
(1) Fault detection
(2) Fault diagnosis (Classification)
(3) Root cause identification
https://github.com/PHME-Datachallenge/Data-Challenge-2021
2021 PHM NA
Turbofan Engine
(1) RUL prediction in a fleet of aircraft engines https://data.phmsociety.org/2021-phm-conference-data-challenge/
2022 PHM EU
Printed Circuit Boards
(1) Prediction of defects in AOI based on SPI
(2) Prediction of human inspection label
(3) Prediction of human-assigned repair label
https://github.com/PHME-Datachallenge/Data-Challenge-2022
2022 PHM NA
Rock Drill
(1) Multiclass fault classification for rock drills https://data.phmsociety.org/2022-phm-conference-data-challenge/
2023 PHM AP
Spacecraft Propulsion System
(1) Diagnose normal, bubble anomalies
(2) Diagnose solenoid valve faults
(3) Diagnose unknown abnormal cases
https://phmap.jp/program-data/
2023 IEEE
Gearbox
(1) Fault diagnosis (Classification)
http://www.ieeereliability.com/PHM2023/index.html
or http://www.ieeereliability.com/PHM2023/assets/data-challenge.pdf